OpenStudy (tux):

Decide convergence / divergence: sum n=1 to infinity 1/(n*sqrt(n)+2)

OpenStudy (anonymous):

converges for sure

OpenStudy (anonymous):

since exponent in denominator is greater than 1

OpenStudy (tux):

How to solve using comparison test or integral test?

OpenStudy (amistre64):

integral test is just integrate from 1 to infinity

OpenStudy (amistre64):

comparison is a limit of the ratio of .... \[\frac{a_{n+1}}{a_n}\]

OpenStudy (anonymous):

actually that is "ratio test" comparison means compare to something else you know converges (or diverges)

OpenStudy (amistre64):

of is the comparison a b_n of another ..... yeah lol

OpenStudy (anonymous):

\[\sum_1^{\infty}\frac{1}{n\sqrt{n}+2}\]compare \[\frac{1}{n\sqrt{n}+2}<\frac{1}{n^{\frac{3}{2}}}\] second sum converges